Calculate inflation rate from 2016 to 2017 when the index number in 2016 is 137 and...

  1. Calculate inflation rate from 2016 to 2017 when the index number in 2016 is 137 and in 2017 is 140.

  2. Calculate the GDP loss if the unemployment rate in the country is 7.9% and actual GDP processed is $1,090.

  3. What would be the labor force participation rate and the unemployment rate in the country if the adult population is 190 million with 165 working and 25 million looking for jobs?
Answer #1

1. Inflation rate from 2016 to 2017

= [(140 - 137) / 137] * 100 = 2.19%

3. Here, 165 people are employed. 25 people are looking for job means they are unemployed. So,

Labor force = employed + unemployed = 165 + 25 = 190

Labor force participation rate

= (Labor force / Adult population) * 100

= (190 / 190) * 100

= 100%

Unemployment rate = (Unemployed / Labor force) * 100

= (25 / 190) * 100

= 13.16%